Please use Excel and include all formulas in different cells so that it is clearH A B C D E F G 1 Northeast Hospital is analyzing a potential project for a new outpatient center Please use the following fa
Answer #1

Based upon the assumptions the projected numbers are calculated as shown below:

F3 D E F Н - J K L M N a 7 8 52500 Year o 9 $ $ 10 Total projected visits Average Revenue per visit Average Variable costs pe

N47 fx D E F G H J K L M N 0 26 27 Year o 28 $ Year 1 Year 2 1,500 $ 93,675 $ 7,50,000 $ 7,50,000 $ Year 3 1,52,456 s 7,50,00

Formulas are calculated as below:

N47 G H J K L M N 7 8 Year o Year 1 ear 1 9 =F8 =F9 10 Year 2 = 19*(1+10%) =J10*(1+2%) =J11*(1+2%) =)12*(1+1%) =J13*(1+2.5%)

Function Library Derined Names Formula Auditing Calcu N48 X - J K L N Year 5 27 Year o Year 1 Year 2 Year 3 M Year 4 =M25 =M2